OpenStudy (laronjaxon):

can someone help me please :/ its ^11√x5 * √x (simplify)

jhonyy9 (jhonyy9):

so this you can writing using exponent form like x^(5/11) * x^(1/2) =x^(5/11 +1/2) =x^((10+11)/22) =x^(21/22)

jhonyy9 (jhonyy9):

ok ?

OpenStudy (laronjaxon):

thank you so much! so the answer would be ^22√x21? right?

jhonyy9 (jhonyy9):

yes sure

OpenStudy (laronjaxon):

thank youu!

jhonyy9 (jhonyy9):

yw good luck bye